# Cosmic Ray Magnetohydrodynamics: A New Two-Moment Framework with Numerical Implementation

### v1

@@ Abstract
Cosmic rays (CRs) play a pivotal role in various astrophysical systems, delivering feedback over a broad range of scales. However, modeling CR transport remains challenging due to its inherently multi-scale nature and complex microphysics. Recent advances in two-moment CR hydrodynamics have alleviated some of these challenges, improving understanding of CR feedback. Yet, current two-moment methods may not be able to directly incorporate all relevant CR transport processes, while the outcome of CR feedback sensitively depends on these underlying microphysics. Furthermore, numerical challenges persist, including instabilities from streaming terms and ambiguities in solver design for coupled CR-MHD systems. In this work, we develop a two-moment description for CR hydrodynamics from first principles. Beyond canonical CR streaming, our formulation accounts for CR pressure anisotropy and Alfvén waves propagating in both directions along the magnetic field, providing a general framework to incorporate more CR transport physics. We implement this framework as a new CR fluid module in the Athena++ code, and validate it through a suite of benchmark tests. In particular, we derive the full dispersion relation of the two-moment CR-MHD system, identifying the CR-acoustic instability as well as other wave branches. These CR-MHD waves serve as rigorous benchmarks and also enable the use of realistic signal speeds in our Riemann solver. We propose a time step guideline to mitigate numerical instabilities arising from streaming source terms.
1 Introduction
@@
Cosmic rays (CRs) are high-energy charged particles traversing the Universe at velocities close to the speed of light. Despite their relativistic speeds, CRs do not freely escape from the galactic disk at the light-crossing time scale, but are instead confined for a time scale $\sim 10^{4}$ times longer, leading to CR energy densities roughly in equipartition with other components of the interstellar medium (ISM) including thermal, turbulent, radiation and magnetic energy densities in galaxies (Boulares & Cox, 1990; Ferrière, 2001; Elmegreen & Scalo, 2004; Cox, 2005; Grenier et al., 2015). Effects of CRs are thus considered as an integral part of ISM dynamics (see Zweibel, 2017; Ruszkowski & Pfrommer, 2023, for reviews). Recent advances in observational and computational techniques have renewed interest in CRs due to their known or potential roles in ionizing gas (Padovani et al., 2009), suppressing star formation (e.g., Jubelgas et al., 2008; Chen et al., 2016; Farcy et al., 2022), launching galactic winds (for very recent work, see e.g., Sike et al., 2025; Thomas et al., 2025, additional earlier studies are listed below), modifying the phase structure in the circumgalactic medium (CGM) (e.g., Sharma et al., 2010b; Ji et al., 2020; Huang et al., 2022; Weber et al., 2025), and modulating black hole feedback (e.g., Sijacki et al., 2008; Guo & Oh, 2008; Lin et al., 2023; Su et al., 2025). It has become essential for studies of galaxy formation and evolution to incorporate CR feedback properly, as CRs are actively involved in a vast array of processes in the life cycle of a galactic ecosystem.
@@
Modeling CR feedback numerically, however, remains challenging due to its multi-scale nature. For $\sim\rm GeV$ CRs, which dominate the CR energy distribution and are dynamically the most important, gyroradii are typically $\lesssim 1$AU (or $\sim 10^{-6}$pc). Directly modeling both microphysical kinetics and galactic scales is prohibitive, given the dynamic range spanning 10 orders of magnitude (Zweibel, 2017). To study large scale one therefore must coarse-grain the kinetic physics of CRs, and treat CR ensembles as a bulk fluid. The fluid formulation for CRs was initially developed by integrating over momentum the Fokker-Plank equation that describes CR transport (e.g., as in Skilling, 1971, 1975; Schlickeiser, 2002; Zank, 2014). Assuming a nearly isotropic CR distribution, an energy-weighted integration leads to a fluid equation governing the evolution of CR energy density (e.g., as in Drury & Voelk, 1981; McKenzie & Voelk, 1982; Voelk et al., 1984; Breitschwerdt et al., 1991; Pfrommer et al., 2017):
@@
Here $\mathcal{E}_{\rm cr}$, $P_{\rm cr}\equiv\mathcal{E}_{\rm cr}/3$ and $\kappa$ represent CR energy density, CR pressure, and the CR diffusion coefficient, respectively, $\vec{u}$ denotes the MHD gas velocity and $\vec{b}\equiv\vec{B}/|\vec{B}|$ is the unit vector along the large-scale background magnetic field. $\vec{v}_{A}\equiv\vec{B}/(4\pi\rho)^{1/2}$ is the traditional Alfvén velocity and $\vec{v}_{\rm st}\equiv-{\rm sgn}(\vec{b}\cdot\nabla P_{\rm cr})\vec{v}_{A}$ is the streaming velocity, which points down the gradients in the CR pressure.
@@
Coupling the one-moment CR hydrodynamic evolution equation, Equation 1, with MHD equations enabled early CR feedback models that demonstrated the possibility of CRs driving galactic winds (e.g., Ipavich, 1975; Breitschwerdt et al., 1991; Zirakashvili et al., 1996; Samui et al., 2010; Dorfi & Breitschwerdt, 2012; Recchia et al., 2017; Mao & Ostriker, 2018; Quataert et al., 2022b, a). However, this one-moment equation is known to induce numerical instabilities through the streaming velocity term $\vec{v}_{\rm st}$, which is ill-defined and flips sign at $P_{\rm cr}$ extrema. To address this step-function-like discontinuity, Sharma et al. (2010a) add a numerical diffusion term to regularize the sharp transition, but at the cost of involving more computational expense and artificial diffusion.
@@
Drawing parallels with radiation hydrodynamics (RHD), Jiang & Oh (2018) developed an alternative two-moment framework that additionally evolves a separate equation for CR energy flux (see also Chan et al. (2019) for a similar implementation). Independently, Thomas & Pfrommer (2019) derived a similar two-moment formalism from first principles base on a moment hierarchy of the CR Fokker-Planck equation. They additionally evolve a wave subsystem to account for the coupling between CRs and the MHD gas. By incorporating a reduced speed of light to accelerate integration (as developed for the RHD context by, e.g. Gnedin & Abel, 2001; Skinner & Ostriker, 2013), two-moment methods retain computational tractability and have become the new norm in CR feedback modeling. Despite their significant advantages over previous numerical approaches, two-moment methods for CR hydrodynamics still have limitations. Perhaps chief among these are the uncertainties in the microphysical CR transport model, and the challenges of integrating kinetic effects into fluid-scale descriptions to accurately reflect macroscopic energy and momentum transfer.
@@
Due to their low number densities and relativistic speeds, GeV CRs are nearly collisionless particles, minimally interacting with the background ions, atoms, and electrons. Instead, they exchange energy and momentum efficiently with the thermal plasmas through resonant scattering by small-scale electromagnetic fluctuations, whose wavelengths are comparable to the CR gyro-radii (Ruszkowski & Pfrommer, 2023). Origins of such magnetic fluctuations include cascade of large-scale MHD turbulence and MHD waves driven by plasma instabilities. While high-energy CRs $(\gtrsim 100{\rm GeV})$ are largely confined by external turbulence (e.g., Chandran, 2000; Yan & Lazarian, 2011), lower-energy CRs—most relevant to ISM dynamics and chemistry—are primarily scattered by MHD waves generated by CRs themselves via gyro-resonant instabilities (thus called “self-confinement”, e.g., Kulsrud & Pearce, 1969; Wentzel, 1974), since the external turbulence cascades to very low amplitudes at the small resonant scales of low-energy CRs (Blasi et al., 2012; Evoli et al., 2018). In the self-confinement scenario, the most widely studied and often dominant gyro-resonant instability is the cosmic-ray streaming instability (CRSI): large-scale CR pressure gradients drive collective CR bulk motion (i.e., streaming) in the direction $\vec{b}\vec{b}\cdot\nabla P_{\rm cr}$, which excites Alfvén waves when the drift velocity exceeds $v_{A}$. Within the reference frame of waves, the CRs are resonantly scattered towards isotropy, effectively locking their bulk flow to the local Alfvén speed if the wave amplitude (as limited by damping) is sufficiently high. By scattering CRs and damping via interactions with the background plasma, MHD waves facilitate energy and momentum transfer between CRs and the MHD gas. Growth and saturation of CRSI, including wave damping, has been extensively studied via MHD-PIC simulations (Bai et al., 2019; Bai, 2022; Plotnikov et al., 2021; Bambic et al., 2021)
@@
To test the paradigm of galactic-scale CR self-confinement with CRSI, Armillotta et al. (2021) implemented a subgrid model for CR scattering with amplitudes set by locally balancing CRSI wave excitation with ion-neutral and nonlinear Landau wave damping (see also Hopkins et al., 2021; Sike et al., 2025; Thomas et al., 2025, for similar implementations), in combination with the two-moment CR solver of Jiang & Oh (2018). This was applied to study transport of GeV CR protons in realistic multiphase gas and magnetic fields, as obtained from high-resolution star-forming ISM simulations for both solar neighborhood and other environmental conditions (Armillotta et al., 2022). Applying an extension of the model that allows for multiple energy groups and CR electrons, Linzer et al. (2025); Armillotta et al. (2025) showed that both CR proton and electron spectra agree with observations at 11001-100GeV, and energy-dependent grammage also satisfies observed constraints. This initial success in combining an instability-driven scattering subgrid model with a two-moment solver to investigate CR feedback effects is highly encouraging, and motivates efforts to refine and extend CR-MHD methods to represent gyrokinetic effects beyond CRSI.
@@
Notably, the CR pressure anisotropy instability (CRPAI) has been identified as a significant alternative process that drives CR self-confinement (first described by Kulsrud (2004) and further studied by Lazarian & Beresnyak (2006); Yan & Lazarian (2011); Lebiga et al. (2018); Zweibel (2020); Sun et al. (2024)). In this scenario, an initially isotropic CR distribution in phase space becomes prolate/oblate due to adiabatic invariance as CR fluid elements experience rapidly increasing/decreasing magnetic fields. The resulting CR pressure anisotropy then supplies free energy for wave growth. Distinct from the CRSI, the CRPAI generates both forward and backward propagating Alfvén waves and with no preference for the $\nabla P_{\rm cr}$ direction. This characteristic makes CRPAI an important complement to the CRSI, especially at the confounding CR bottleneck structures where the CR and thermal pressure gradients are misaligned, flattening $\nabla P_{\rm cr}$ and deactivating the CRSI (Wiener et al., 2017). Consequently, the additional coupling provided by CRPAI could influence CR-cloud interactions (e.g., Wiener et al., 2017; Brüggen & Scannapieco, 2020) and regulate wind launching and mass loading in CR-overpressured regions (e.g., bubbles, see Schroer et al., 2022; Lancaster et al., 2024). Also, CRPAI itself might dominate the wave excitations in environments with temporally or spatially rapidly changing magnetic fields, providing near-source heating (e.g., Wagner et al., 2007; Caprioli et al., 2009), modifying mixing layers in the multiphase ISM/CGM (Ji et al., 2019; Fielding et al., 2023; Zhao & Bai, 2023) and affecting the thermal and hydrostatic stability of CGM halos (e.g., Tsung et al., 2022).
@@
As uncertainties in CR feedback are intrinsically tied to unresolved transport physics, advancing CR feedback modeling to to next level requires a general framework capable of incorporating all potential transport mechanisms. However, current two-moment schemes have limited extensibility. The formulation in Jiang & Oh (2018) implicitly assumes that Alfvén waves can only propagate along one direction (either along or opposed to the local $\vec{b}$). This is well motivated by CRSI, but is not consistent with more general Alfvén wave configurations as in CRPAI. The Thomas & Pfrommer (2019) formulation in principle does allow general wave configurations, but their derivation only targets drift anisotropy, and incorporating additional physics would lead to further complications, especially in dealing with the wave subsystem.
@@
In this work, we take a first-principles starting point similar to Thomas & Pfrommer (2019), and formulate a two-moment description for CR hydrodynamics that accommodates an arbitrary configuration of Alfvén waves interacting with the CR ensemble. We include CR pressure anisotropy along with its effects in both the scattering term and the advection term. Our formulation provides a more general framework for CR feedback modeling, which has the flexibility to incorporate various CR transport modes such as the extrinsic turbulence, self-confinement by the CRSI, the CRPAI and potentially their combination. We implement our two-moment CR formulation into the Athena++ code as a fluid species coupled to the MHD gas. To ensure numerical robustness, we also calculate characteristic wave speeds for the CR-MHD system and apply them in our HLLE Riemann solver. Furthermore, we perform von-Neumann stability analysis and derive a time-step criterion.
@@
This paper is organized as follows: Section 2 derives the framework for our CR hydrodynamics from first principles, beginning with the CR kinetic equation. Section 3 outlines the numerical implementation of our CR module in the Athena++ code, detailing considerations such as wave speeds adopted in the HLLE solver, and time step criterion prescribed to ensure numerical stability. Section 4 and Section 5 present benchmark tests validating code performance. Section 6 discusses our formulation in comparison with prior two-moment schemes, potential astrophysical applications, limitations as well as future extensions of our current implementation. A preliminary analysis of characteristic waves in the CR-MHD system is presented in Appendix A, followed by a von Neumann analysis to obtain the stability condition in Appendix B. In Appendix C, we provide a simple analytical derivation of CR energy density evolution in cylindrical coordinates with CR scattering solely subject to CRPAI, which serves as a benchmark for comparison with our numerical tests.
2 CR hydrodynamics

### v2

@@ Abstract
Cosmic rays (CRs) play a pivotal role in various astrophysical systems, delivering feedback over a broad range of scales. However, modeling CR transport remains challenging due to its inherently multi-scale nature and complex microphysics. Recent advances in two-moment CR hydrodynamics have alleviated some of these challenges, improving our understanding of CR feedback. Yet, current two-moment methods may not be able to directly incorporate all relevant CR transport processes, while the outcome of CR feedback sensitively depends on these underlying microphysics. Furthermore, numerical challenges persist, including instabilities from streaming terms and ambiguities in solver design for coupled CR-MHD systems. In this work, we develop a two-moment description for CR hydrodynamics from first principles. Beyond canonical CR streaming, our formulation accounts for CR pressure anisotropy and Alfvén waves propagating in both directions along the magnetic field, providing a general framework to incorporate more CR transport physics. We implement this framework as a new CR fluid module in the Athena++ code, and validate it through a suite of benchmark tests. In particular, we derive the full dispersion relation of the two-moment CR-MHD system, identifying the CR-acoustic instability as well as other wave branches. These CR-MHD waves serve as rigorous benchmarks and also enable the use of realistic signal speeds in our Riemann solver. We propose a time step guideline to mitigate numerical instabilities arising from streaming source terms.
I Introduction
@@
Cosmic rays (CRs) are high-energy charged particles traversing the Universe at velocities close to the speed of light. Despite their relativistic speeds, CRs do not freely escape from the galactic disk at the light-crossing time scale, but are instead confined for a time scale $\sim 10^{4}$ times longer, leading to CR energy densities roughly in equipartition with other components of the interstellar medium (ISM), including thermal, turbulent, radiation and magnetic energy densities in galaxies (Boulares and Cox, 1990; Ferrière, 2001; Elmegreen and Scalo, 2004; Cox, 2005; Grenier et al., 2015). Effects of CRs are thus considered as an integral part of ISM dynamics (see Zweibel, 2017; Ruszkowski and Pfrommer, 2023; Hopkins, 2025, for reviews). Recent advances in observational and computational techniques have renewed interest in CRs due to their known or potential roles in ionizing gas (Padovani et al., 2009), regulating star formation (e.g., Jubelgas et al., 2008; Chen et al., 2016; Farcy et al., 2022), driving galactic winds (for very recent work, see e.g., Sike et al., 2025; Thomas et al., 2025a, additional earlier studies are listed below), shaping the phase structure in the circumgalactic medium (CGM) (e.g., Sharma et al., 2010b; Ji et al., 2020; Huang et al., 2022; Weber et al., 2025), and modulating black hole feedback (e.g., Sijacki et al., 2008; Guo and Oh, 2008; Lin et al., 2023; Su et al., 2025). It has become essential for studies of galaxy formation and evolution to incorporate CR feedback properly, as CRs are actively involved in a vast array of processes in the life cycle of a galactic ecosystem.
@@
Modeling CR feedback numerically, however, remains challenging due to its multi-scale nature. For $\sim\rm GeV$ CRs, which dominate the CR energy distribution and are dynamically the most important, gyroradii are typically $\lesssim 1$AU (or $\sim 10^{-6}$pc). Directly modeling both microphysical kinetics and galactic scales is prohibitive, given the dynamic range spanning 10 orders of magnitude (Zweibel, 2017). To study large scale one therefore must coarse-grain the kinetic physics of CRs, and treat CR ensembles as a bulk fluid. The fluid formulation for CRs was initially developed by integrating over momentum the Fokker-Plank equation that describes CR transport (e.g., as in Skilling, 1971, 1975; Schlickeiser, 2002; Zank, 2014). Assuming a nearly isotropic CR distribution, an energy-weighted integration leads to a fluid equation governing the evolution of CR energy density (e.g., as in Drury and Voelk, 1981; McKenzie and Voelk, 1982; Voelk et al., 1984; Breitschwerdt et al., 1991; Pfrommer et al., 2017):
@@
Here $\mathcal{E}_{\rm cr}$, $P_{\rm cr}\equiv\mathcal{E}_{\rm cr}/3$ and $\kappa$ represent CR energy density, CR pressure, and the CR diffusion coefficient, respectively. $\vec{u}$ denotes the MHD gas velocity and $\vec{b}\equiv\vec{B}/|\vec{B}|$ is the unit vector along the large-scale background magnetic field. $\vec{v}_{A}\equiv\vec{B}/(4\pi\rho)^{1/2}$ is the traditional Alfvén velocity and $\vec{v}_{\rm st}\equiv-{\rm sgn}(\vec{b}\cdot\nabla P_{\rm cr})\vec{v}_{A}$ is the streaming velocity, which points down the gradients in the CR pressure.
@@
Coupling the one-moment CR hydrodynamic equation, Equation 1, with MHD equations enabled early CR feedback models that demonstrated the possibility of CRs driving galactic winds (e.g., Ipavich, 1975; Breitschwerdt et al., 1991; Zirakashvili et al., 1996; Samui et al., 2010; Dorfi and Breitschwerdt, 2012; Recchia et al., 2017; Mao and Ostriker, 2018; Quataert et al., 2022b, a). However, this one-moment equation is known to induce numerical instabilities through the streaming velocity term $\vec{v}_{\rm st}$, which is ill-defined and flips sign at $P_{\rm cr}$ extrema. To address this step-function-like discontinuity, Sharma et al. (2010a) added a numerical diffusion term to regularize the sharp transition, but at the cost of involving more computational expense and artificial diffusion.
@@
Drawing parallels with radiation hydrodynamics (RHD), Jiang and Oh (2018) developed an alternative two-moment framework that additionally evolves a separate equation for CR energy flux (see also Chan et al. (2019) for a similar implementation). Independently, Thomas and Pfrommer (2019) derived a similar two-moment formalism from first principles based on a moment hierarchy of the CR Fokker-Planck equation. They additionally evolve a wave subsystem to account for the coupling between CRs and the MHD gas. By incorporating a reduced speed of light to accelerate integration (as developed for the RHD context by, e.g. Gnedin and Abel, 2001; Skinner and Ostriker, 2013), two-moment methods retain computational tractability and have become the new norm in CR feedback modeling. Despite their significant advantages over previous numerical approaches, two-moment methods for CR hydrodynamics still have limitations. Perhaps chief among these are the uncertainties in the microphysical CR transport model, and the challenges of integrating kinetic effects into fluid-scale descriptions to accurately reflect macroscopic energy and momentum transfer.
@@
Due to their low number densities and relativistic speeds, GeV CRs are nearly collisionless particles, minimally interacting with the background ions, atoms, and electrons. Instead, they exchange energy and momentum efficiently with the thermal plasmas through resonant scattering by small-scale electromagnetic fluctuations, whose wavelengths are comparable to the CR gyro-radii (Ruszkowski and Pfrommer, 2023). Origins of such magnetic fluctuations include cascade of large-scale MHD turbulence and MHD waves driven by plasma instabilities. While high-energy CRs $(\gtrsim 100{\rm GeV})$ are largely confined by external turbulence (e.g., Chandran, 2000; Yan and Lazarian, 2011), lower-energy CRs—most relevant to ISM dynamics and chemistry—are primarily scattered by MHD waves generated by CRs themselves via gyro-resonant instabilities (thus called “self-confinement”, e.g., Kulsrud and Pearce, 1969; Wentzel, 1974), since the external turbulence cascades to very low amplitudes at the small resonant scales of low-energy CRs (Blasi et al., 2012; Evoli et al., 2018). In the self-confinement scenario, the most widely studied and often dominant gyro-resonant instability is the CR streaming instability (CRSI): large-scale CR pressure gradients drive collective CR bulk motion (i.e., streaming) in the direction $\vec{b}\vec{b}\cdot\nabla P_{\rm cr}$, which excites Alfvén waves when the drift velocity exceeds $v_{A}$. Within the reference frame of waves, the CRs are resonantly scattered towards isotropy, effectively locking their bulk flow to the local Alfvén speed if the wave amplitude (as limited by damping) is sufficiently high. By scattering CRs and damping with the background plasma, MHD waves facilitate energy and momentum transfer between CRs and the MHD gas. MHD-PIC simulations have extensively investigated the growth of CRSI, the role of wave damping, and the resulting saturation of the instability (Bai et al., 2019; Bai, 2022; Plotnikov et al., 2021; Bambic et al., 2021; Lemmerz et al., 2024).
@@
To test the paradigm of galactic-scale CR self-confinement with CRSI, Armillotta et al. (2021) implemented a subgrid model for CR scattering, in which amplitudes are set by locally balancing CRSI wave excitation with ion-neutral and nonlinear Landau wave damping (see also Hopkins et al., 2021; Sike et al., 2025; Thomas et al., 2025a, for similar implementations), coupled with the two-moment CR solver of Jiang and Oh (2018). This framework was applied to study transport of GeV CR protons in realistic multiphase gas and magnetic fields, as obtained from high-resolution star-forming ISM simulations for both solar neighborhood and other environmental conditions (Armillotta et al., 2022). Extending the model to multiple energy groups and including CR electrons, Linzer et al. (2025); Armillotta et al. (2025) showed that both CR proton and electron spectra match observations across 11001-100GeV, with energy-dependent grammage also satisfying observational constraints. This initial success highlights the promise of two-moment frameworks equipped with subgrid CR transport, and motivates advanced CR–MHD methods that include gyrokinetic effects beyond CRSI along with more robust numerical schemes.
@@
Notably, the CR pressure anisotropy instability (CRPAI) has been identified as a significant alternative process that drives CR self-confinement (first described by Kulsrud (2004) and further studied by Lazarian and Beresnyak (2006); Yan and Lazarian (2011); Lebiga et al. (2018); Zweibel (2020); Sun et al. (2024)). In this scenario, an initially isotropic CR distribution in phase space becomes prolate/oblate due to adiabatic invariance as CR fluid elements experience rapidly increasing/decreasing magnetic fields. The resulting CR pressure anisotropy then supplies free energy for wave growth. Distinct from the CRSI, the CRPAI generates both forward and backward propagating Alfvén waves with no preference for the $\nabla P_{\rm cr}$ direction. This characteristic makes CRPAI an important complement to the CRSI, especially at the confounding CR bottleneck structures where the CR and thermal pressure gradients are misaligned, flattening $\nabla P_{\rm cr}$ and deactivating the CRSI (Wiener et al., 2017). Consequently, the additional confinement provided by CRPAI could influence CR-cloud interactions (e.g., Wiener et al., 2017; Brüggen and Scannapieco, 2020) and regulate wind launching and mass loading in CR-overpressured regions (e.g., bubbles, see Schroer et al., 2022; Lancaster et al., 2024). Also, CRPAI itself might dominate the wave excitations in environments with temporally or spatially rapidly changing magnetic fields, providing near-source heating (e.g., Wagner et al., 2007; Caprioli et al., 2009), modifying mixing layers in the multiphase ISM/CGM (Ji et al., 2019; Fielding et al., 2023; Zhao and Bai, 2023) and affecting the thermal and hydrostatic stability of CGM halos (e.g., Tsung et al., 2022).
@@
As uncertainties in CR feedback are intrinsically tied to unresolved transport physics, advancing CR feedback modeling to next level requires a general framework capable of incorporating all potential transport mechanisms. However, current two-moment schemes have limited extensibility. The formulation in Jiang and Oh (2018) implicitly assumes that Alfvén waves can only propagate along one direction (either along or opposed to the local $\vec{b}$). This is well motivated by CRSI, but is not consistent with more general Alfvén wave configurations as in CRPAI. The Thomas and Pfrommer (2019) formulation in principle does allow general wave configurations, but their derivation only targets drift anisotropy, and incorporating additional physics would lead to further complications, especially in dealing with the wave subsystem.
@@
In this work, we take a first-principles starting point similar to Thomas and Pfrommer (2019), and formulate a two-moment description for CR hydrodynamics that accommodates an arbitrary configuration of Alfvén waves interacting with the CR ensemble. We include CR pressure anisotropy along with its effects in both the scattering term and the advection term. Our formulation provides a more general framework for CR feedback modeling, which has the flexibility to incorporate various CR transport modes such as the extrinsic turbulence, self-confinement by the CRSI, the CRPAI and potentially their combination. We implement our two-moment CR formulation into the Athena++ code (Stone et al., 2020) as a fluid species coupled to the MHD gas. To ensure numerical robustness, we also calculate characteristic wave speeds for the CR-MHD system and apply them in our HLLE Riemann solver. Furthermore, we perform von-Neumann stability analysis and derive a time-step criterion.
@@
This paper is organized as follows: Section II derives the framework for our CR hydrodynamics from first principles, beginning with the CR kinetic equation. Section III outlines the numerical implementation of our CR module in the Athena++ code, detailing considerations such as wave speeds adopted in the HLLE solver, and time step criterion prescribed to ensure numerical stability. Section IV and Section V present benchmark tests validating code performance. Section VI discusses our formulation in comparison with prior two-moment schemes, potential astrophysical applications, limitations as well as future extensions of our current implementation. A preliminary analysis of characteristic waves in the CR-MHD system is presented in Appendix A, followed by a von Neumann analysis to obtain the stability condition in Appendix B. In Appendix C, we provide a simple analytical derivation of CR energy density evolution in cylindrical coordinates with CR scattering solely subject to CRPAI, which serves as a benchmark for comparison with our numerical tests.
II CR hydrodynamics